RELAMPAGO servicing events

s9: The solar panels blew down during a high wind 11/12 Jun.  On 14 Jun, the flux sensors were powered down remotely, but the station still died that night.  On 21 Jun, one panel was reconnected to restart the station.  Both panels were reconnected on 18 Jul.

s4: The radiometer tripod was knocked down and soils dug up by cattle on 14 Jun.  On 26 Jun the station went down completely when all usb devices were stolen from the station.  On 24 Jul the station was mostly restored.

s12: On 14 Jul the mote cable was found disconnected and plugged back in.

s13: On 19 Jul, station found down and restarted.  Mote and 2D also found not working, so changed DSM ports.

s10: On 20 Jul, station restarted after dying after visit on 14 Jul.

s15: On 28 Jul, the TRH died.  Did not respond to usual fixes on 19 Oct and was replaced on 21 Oct.

s5: On 7 Sep, the mote was bad.

s12: On 7 Sep, the mote was not reporting.

s10: On 17 Oct, the TRH fan wasn't working, so the TRH was replaced.

s12: On 18 Oct, the soils and radiometer had been damaged by livestock.  The radiometer and Tsoil were restored.

s2: On 19 Oct, the tipping bucket was resetting, but appeared to fix itself on the same day.  Replaced on 11 Nov.

s5: On 8 Nov, TP01 found out of the ground and was replaced.

s6: On 8 Nov, reconnected the Qsoil.

s7: On 10 Nov, replaced Tsoil.

s4: On 10 Nov, cleared out clogs in tipping bucket.

s14: On 12 Nov, PTB was bad and was replaced.  Qsoil and TP01 were disconnected and were completely replaced.

s9: On 14 Nov, Tsoil was found to be exposed, so reburied.

s12: On 14 Nov, tipping bucket was clogged, so cleaned.  Mote was rebooting, so replaced.  Qsoil, Gsoil, TP01 were removed by livestock, so replaced. Rad stand moved a bit.  On 15 Nov, changed TP01 wisard board to get it going, so calibration off after that.

s11: On 14 Nov, tipping bucket was clogged, so cleaned.  Soil and radiometer cables unplugged, but working once reconnected.

s1: On 16 Nov, mote rebooting.

s2: On 16 Nov, Qsoil not reporting.  Found chewed and replaced on 18 Nov.

s4: On 16 Nov, TP01 reporting bad.

s9: On 16 Nov, Tcase is high -- perhaps bad connection internally.  Logbook comment that day suggests a fix.  NR01 replaced 17 Nov.

s7: On 17 Nov, mote cable reconnected.

s6: On 17 Nov, many mote cables were reconnected and mote replaced due to bad port.

s4: On 22 Nov, NR01 died. Reconnected 13 Jan.

s5: On 29 Nov, mote mostly dead.  Cables reconnected 19 Dec, then cable replaced 20 Jan.

s7: On 29 Nov, TP01 found mostly dead since Oct.  Tsoil bad since 18 Dec.  Qsoil bad since 7 Dec.  All fixed on 15 Jan.

s6: On 14 Dec, Tsoil reading 0.  Replaced 14 Jan.

s9: On 9 Jan, Tsoil and TP01 not reporting.

s13: On 30 Dec, NR01 stopped reporting.  Replaced and tripod releveled 17 Jan.

s9: On 24 Nov, mote mostly bad.  Died completely on 28 Nov.  Fixed by disconnecting NR01 on 19 Dec.  Still not reconnected on 25 Dec. 17 Jan mote replaced, Gsoil, TP01, NR01 back up.

s15: On 27 Nov, mote mostly bad.  Sensors reconnected 18 Dec.

s12: On 27 Nov, suspect that Gsoil is upside down? (have later found some Gsoils wired backwards).

s1: On 14 Dec, TRH dead.  Fixed by remote power cycle on 10 Jan.

s3: On 21 Dec, TP01 dead. Replaced 12 Jan.

s9: On 25 Dec, Qsoil reporting low.

s14: On 10 Jan, Tsoil bad.  Mote rebooting intermittently. Tsoil replaced 17 Jan.

s15: On 11 Jan, NR01 releveled.

s1: 11 Jan, tipping bucket clogged

s3: 12 Jan, tipping clogged.

s5: 14 Jan, tipping clogged.

s7: 15 Jan, tipping clogged.

s8: 15 Jan, tipping clogged; soil sensors exposed; TRH bad

s10: 15 Jan, tipping clogged.

s13: 17 Jan, reburied Tsoil;

s10: 20 Jan, Tsoil replaced.

s15: 11 Mar, Rad releveled (was off); tipping cleaned

s1: 12 Mar, Rad releveled (was off); TRH housing replaced; tipping cleaned

s2: 12 Mar, tipping cleaned

s14: 13 Mar, tipping cleaned

s9: 14 Mar, tipping clogged

s12: 14 Mar, tipping clogged

s11: 14 Mar, tipping clogged; TRH SHT replaced

s4: 15 Mar, tipping cleaned

s5: 15 Mar, tipping cleaned

s3: 15 Mar, tipping PIC replaced

s8 18 Mar, replaced PTB210 with 220 in different port

s1: 12 Apr, mote not reporting

s5: 12 Apr, TP01 not reporting

s7: 12 Apr, Tsoil all negative

s8: 12 Apr, all USB devices down

s9: 12 Apr, TP01 odd

s11: 12 Apr, all USB sensors down

s14: 12 Apr, Tsoil not reporting